### v1

# Assessing the Limitations of Large Language Models in Clinical Fact Decomposition

Abstract
Verifying factual claims is critical for using large language models (LLMs) in healthcare. Recent work has proposed fact decomposition, which uses LLMs to rewrite source text into concise sentences conveying a single piece of information, as an approach for fine-grained fact verification. Clinical documentation poses unique challenges for fact decomposition due to dense terminology and diverse note types. To explore these challenges, we present FactEHR, a dataset consisting of full document fact decompositions for 2,168 clinical notes spanning four types from three hospital systems. Our evaluation, including review by clinicians, highlights significant variability in the quality of fact decomposition for four commonly used LLMs, with some LLMs generating 2.6x more facts per sentence than others. The results underscore the need for better LLM capabilities to support factual verification in clinical text. To facilitate future research in this direction, we plan to release our code at https://github.com/som-shahlab/factehr.
Assessing the Limitations of Large Language Models in
Clinical Fact Decomposition
Monica Munnangi*1,2, Akshay Swaminathan*3, Jason Alan Fries*2, Jenelle Jindal4, Sanjana Narayanan2, Ivan Lopez3, Lucia Tu2, Philip Chung7, Jesutofunmi A. Omiye3,8, Mehr Kashyap3, Nigam Shah5,6,9 1Khoury College of Computer Sciences, Northeastern University; 2Center for Biomedical Informatics Research, Stanford University; 3Department of Biomedical Data Science∗∗; 4Stanford Health Care; 5Department of Medicine∗∗; 6Clinical Excellence Research Center∗∗; 7Department of Anesthesiology, Perioperative & Pain Medicine∗∗; 8Department of Dermatology∗∗; 9Technology and Digital Solutions, Stanford Health Care  munnangi.m@northeastern.edu, {akshaysw, jfries}@stanford.edu
**footnotetext: Equal contribution ****footnotetext: Stanford School of Medicine
1 Introduction
Figure 1: Fact decomposition in clinical text with LLMs requires handling diverse document types and complex medical jargon (colors highlight examples).
Verifying and attributing factual claims is essential for using large language models (LLMs) in healthcare. Evaluation strategies were proposed to measure summarization quality, ensure claims are properly grounded (attributed to specific text in the clinical note) (Bannur et al., 2024), text generated for documentation tasks must be grounded in data present in the patient’s EHR. , and reduce hallucinations and related errors (Zheng et al., 2023). A core component of these strategies is fact decomposition—an approach that paraphrases complex text into multiple, concise statements, each expressing a single piece of information (Fabbri et al., 2022; Chen et al., 2023; Min et al., 2023).
Fact decomposition can be used to verify facts against source documents using QA systems, textual entailment, or other model-based evaluation methods. Textual entailment, in particular, has proven effective for automating factuality assessment (Kamoi et al., 2023; Ru et al., 2024; Xie et al., 2024). Although these verification techniques are well-studied in both general and scientific texts (Wadden et al., 2020; Wright et al., 2022), their application to clinical text remains underexplored. This represents a significant gap, as performance metrics are highly sensitive to the quality of fact decomposition (Wanner et al., 2024).
In healthcare, many documentation tasks require summarizing information from electronic health records (EHRs) (Fleming et al., 2023). EHRs encompass multiple facets of patient care, including various types of clinical documents, tabular data, and a variety of unstructured data types (e.g., medical imaging, waveforms). Extracting, rewriting, or verifying evidence from EHRs is necessary for tasks like summarization (Van Veen et al., 2024; Hegselmann et al., 2024), patient phenotyping (Yang et al., 2024), clinical trial recruitment (Wornow et al., 2024), medical text simplification (Devaraj et al., 2021), and knowledge graph construction (Arsenyan et al., 2023). Just as facts in radiology reports must be grounded to specific pixel data (Bannur et al., 2024), text generated for documentation tasks must be grounded in data present in the patient’s EHR. Evaluating the ability of LLMs to perform this attribution is necessary for the successful use of LLMs in healthcare.
The decomposition and verification of the facts in clinical notes using LLMs presents significant challenges. Clinical notes employ medical terminology, special acronyms, and non-grammatical shorthand (see Figure 1). Clinical observations are often compositional, as in: “lobulation at the apex of the left hemithorax along the mediastinal border is residual of slowly resolving hematoma." For research purposes, clinical notes are frequently stripped of their original markup—like tables, lists, and section headers—thereby increasing ambiguity and parsing difficulties. The de-identification process, which masks patient names and dates, further introduces noise that can confound LLMs. Moreover, medical documentation serves diverse functions: nursing notes offer brief, time-stamped updates on a patient’s physiological status, whereas discharge summaries integrate an entire hospital stay’s details for billing and compliance purposes. Consequently, these documents differ substantially in presentation, length, fact density, and temporal scope.
Existing entailment datasets – which denote whether a piece of text (the hypothesis) logically follows from another piece of text (the premise) – do not capture the diversity of clinical documentation, either because they focus on a single note type (Miura et al., 2021), use artificial data (Romanov and Shivade, 2018a), or have a limited representation of hospital systems. Fact decomposition across different types of clinical documents has not been evaluated. In addition, it is frequently unclear how to evaluate the quality of a decomposition because determining entailment requires medical expertise. Finally, frontier model APIs are often not HIPAA-compliant, hindering evaluations using popular datasets such as MIMIC (Johnson et al., 2016). Given these challenges, we explore the following research questions (RQ’s):
RQ1 — Fact decomposition
How do the length, quality, and similarity of fact decompositions differ across LLMs?
RQ2 — Entailment
How accurately can LLMs determine textual entailment of clinical documents?
RQ3 — Fact verification
How accurate and complete are the fact decompositions of clinical notes created by LLMs?
We present FactEHR, a fact decomposition and textual entailment dataset derived from 2,168 clinical notes across four document types from three hospital systems. Using FactEHR, we contribute the following:
  1. 1.
    For RQ1, we analyze the length and similarity of fact decompositions from four LLMs and conduct a qualitative evaluation with clinical experts.
  2. 2.
    For RQ2, we compare LLMs in their ability to perform textual entailment using standard NLI datasets and expert-labeled entailment pairs from clinical documents. We establish feasibility of using GPT-4o to assess entailment.
  3. 3.
    For RQ3, we quantify the accuracy and completeness of fact decompositions from four LLMs.
Our findings reveal considerable variation in the number and atomicity of fact decompositions across clinical documents and across LLMs, with some LLMs generating 2.6x more facts per sentence than others. This raises questions about the validity of metrics that rely on fact decompositions for evaluating LLMs in healthcare documentation tasks.
2 Methods
2.1 Experimental Setup
We use LLMs to perform fact decomposition—the process of breaking down source text into concise, atomic statements that convey individual pieces of information. We then examine the resulting fact decompositions, comparing their lengths and similarity across different LLMs. Finally, we pair each decomposition with its corresponding source note to form entailment pairs, evaluating how accurately and completely LLMs capture the original information.
Models
We consider four LLMs to perform fact decomposition: GPT-4o (OpenAI, 2024), o1-mini (Singhal et al., 2023b), Gemini-1.5-Flash-002 (et al., 2024), and Llama3-8b-Instruct (Dubey et al., 2024). We will refer to Llama3-8b-Instruct as Llama3 or Llama3-8B and Gemini-1.5-Flash-002 as Gemini-1.5 in the rest of the paper (unless specified otherwise). All models were run using HIPAA-compliant compute environments and APIs. Model generation hyperparameters are included in Appendix C.
Prompts
Adapted from (Min et al., 2023), our prompt includes two in-context examples of clinical note fact decompositions and instructs the LLM to output independent facts as a delimited string (Appendix 11).
2.2 Data Sources & Preprocessing
We sample clinical notes from three de-identified research datasets: MIMIC (MIMIC-III (Johnson et al., 2016) and MIMIC-CXR (Johnson et al., 2019)), from Beth Israel Deaconess Medical Center in Boston, MA; CORAL, from the University of California, San Francisco (UCSF) (Sushil et al., 2024); and MedAlign, from Stanford Health Care (SHC), Palo Alto, CA (Fleming et al., 2023). We randomly sample up to 250 notes per type from each dataset, limiting note lengths to 64–3840 whitespace-delimited tokens, for a total of 2,168 notes as summarized in Table 4 and Table 5 for token length. Additional details are in Appendix A.1.
Figure 2: Example fact decompositions across LLMs. Colors indicate mappings from note (left) to corresponding facts, numbered boxes indicate fact counts, and red text indicates hallucination, red boxes indicated missed facts. All LLMs missed the the comparison study (a prior radiograph) and the note’s final report status.
2.3 Note Types
We consider four clinical note types: (1) Procedure Note, which typically includes the procedures, clinical indications, findings, and follow-up recommendations. (2) Nursing Note, which provides a systematic assessment of a patient’s condition across body systems (e.g., cardiovascular, neurologic) at a specific time, with less emphasis on future care planning. (3) Progress Note, which summarizes a patient’s medical status from the previous day and outline the care plan for the next. (4) Discharge Summary, which provides a concise overview of the patient’s presentation, past medical history, key findings, future medical plans, and discharge medications for subsequent care providers.
2.4 RQ1 — Fact Decomposition
We examine the length and similarity of fact decompositions across LLMs, as well as the atomicity of source sentences. Figure 2 outlines example decompositions across all LLMs used in this study.
Fact Counts
We report the average number of facts generated per source note sentence across models and note types. To assess the variation in the number of facts generated across LLMs, we compute the coefficient of variation (standard deviation divided by the mean) of the number of facts generated by each LLM for each document (Everitt, 2006) of these counts. This metric, calculated across all documents within each note type, quantifies the degree of disagreement among LLMs on the total number of facts per document.
Fact Similarities
To quantify the similarity of fact decompositions across LLMs, we use Earth Mover’s Distance (EMD) (Rubner et al., 1998; Kusner et al., 2015), which calculates the minimum cost to transform one probability distribution into another. Specifically, we embed each fact using ClinicalBERT (Alsentzer et al., 2019) and represent a single decomposition as a matrix of these fact embeddings. When comparing two decompositions, one serves as the source distribution of fact embeddings and the other as the target distribution. We then compute EMD using a cosine cost function and uniform weighting to determine how closely the source’s set of facts aligns with the target’s set, yielding a measure of similarity between the two decompositions.
Sentence Atomicity
To quantify the amount of information in each sentence of a source document, we estimate the number of facts contained within a single sentence, which we refer to as sentence atomicity. We assign each fact to its most similar sentence using a greedy matching approach based on cosine similarity of ClinicalBERT embeddings.
2.5 RQ2 — Entailment
To evaluate the accuracy and completeness of fact decompositions, we pair each decomposition with its corresponding source note to form entailment pairs. We investigate whether LLMs can accurately evaluate textual entailment in this setting.
We first benchmark five LLMs—GPT-4o, GPT-4o-mini, Gemini-1.5, Llama3-8B, and Llama3-70B —using four publicly available natural language inference (NLI) benchmarks—SciTail (Khot et al., 2018), MedNLI (Romanov and Shivade, 2018a), MultiNLI (Williams et al., 2018a), and SNLI (Bowman et al., 2015) (Appendix A.2)—and the FactEHR development set (described below).
We tune our prompt using 40 premise-hypothesis pairs sampled from these datasets, optimizing for F1 score. The final prompt, adapted from (Xie et al., 2024), instructs the LLM to output a binary entailment judgement in JSON format (Appendix 12).
We then selected the model with the highest F1 score and applied it to all 987,266 entailment pairs in the FactEHR dataset (details provided below). Of these pairs, 1,036 were manually annotated by clinical experts, allowing us to directly assess the LLM’s accuracy on the FactEHR dataset.
2.6 RQ3 — Fact Verification
To assess the accuracy and completeness of model-generated fact decompositions, we use fact-precision and fact-recall, entailment-based metrics proposed by (Xie et al., 2024).
Let $S$ denote the set of sentences obtained by tokenizing the clinical note $d$, and let $C$ represent the set of factual claims produced by decomposing $d$. We define $[d\models c]=1$ if a hypothesis $c$ is completely entailed by a premise $d$, and 00 otherwise. For example, the hypothesis, “There is a pleural effusion on the left side,” is completely entailed by the premise “Left pleural effusion increased from prior scan.”
Fact Precision
Fact precision measures the accuracy of decomposed claims using entailment, treating each fact as a hypothesis and the entire clinical note as the premise. It is defined as the proportion of facts in $C$ that are entailed by clinical note $d$: $P^{*}=\frac{1}{|C|}\sum_{c}[d\models c]$
Fact Recall
Fact recall measures the completeness of a fact decomposition using entailment, treating each sentence of the clinical note as a hypothesis and the entire set of facts as the premise. It is defined as the proportion of sentences in $S$ that are entailed by the fact decomposition $C$:
$R{\text{*}}=\frac{1}{|S|}\sum_{s\in S}[C\models s],$
where $S$ is the set of sentences in the clinical note, and $[C\models s]$ is an indicator function that evaluates to 1 if $s$ is entailed by $C$, and 0 otherwise.
We use a weighted fact recall metric to account for multiple facts within a single sentence and parsing issues. Instead of counting each sentence as a single fact, we estimate the number of facts per sentence. Sentences that contain incomplete or invalid information (e.g., just units, numbers, or parsing errors) receive a weight of zero. All other sentences are weighted by their estimated fact count, determined by the sentence atomicity measure described earlier. The weighted fact recall is then defined as:
$R{\text{*}}_{w}=\frac{\sum_{s\in S}w_{s}[C\models s]}{\sum_{s\in S}w_{s}},$
where $w_{s}$ is the weight assigned to sentence $s$, representing the estimated number of facts it contains.
Figure 3: Mean EMD (cosine cost) across LLM fact decompositions and source sentences. This measures how embedded fact decompositions vary across LLMs and in comparison to the original document.
3 Results
We introduce FactEHR (pronounced “factor”), a dataset for fact decomposition and textual entailment created from de-identified clinical notes originating from three hospital systems. FactEHR includes:
  • Fact Decompositions: 8,665 fact decompositions produced by four LLMs from 2,168 clinical notes spanning four document types.
  • Entailment Pairs: 987,266 entailment pairs generated by pairing fact decompositions with their source text. Each pair has a binary entailment label from GPT-4o, and a subset of 1,036 pairs also includes labels from clinical experts.
  • Development Set: A separate set of 2,468 entailment pairs annotated by clinical experts, used for entailment model development and tuning.
Sampling
FactEHR includes 8,665 fact decompositions generated by four LLMs from 2,168 clinical documents (4 x 2,168 = 8,172; five decompositions were not produced due to content moderation policies). To evaluate fact precision and recall, we formed 987,266 pairs from these decompositions. Each pair is classified as either a fact-precision pair, where the premise is the clinical note and the hypothesis is a fact from the decomposition, or a fact-recall pair, where the premise is the fact decomposition and the hypothesis is a sentence from the corresponding clinical note.
Entailment annotation
From these 987,266, we randomly selected 1,036 pairs for manual entailment annotation by clinical experts. Clinical experts included two board certified physicians, two residents, two medical students, and a clinical researcher. Annotators were provided instructions for labeling entailment (Appendix D), and a subset of 100 entailment pairs were labeled in duplicate to calculate inter-rater agreement.
Note Type Model Name Mean $\pm$ SE
Procedure Note GPT-4o 1.45 $\pm$ 0.03
o1-mini 1.45 $\pm$ 0.03
Gemini-1.5 1.12 $\pm$ 0.03
Llama3 8B 0.89 $\pm$ 0.03
Nursing Note GPT-4o 2.37 $\pm$ 0.14
o1-mini 2.42 $\pm$ 0.13
Gemini-1.5 2.05 $\pm$ 0.12
Llama3 8B 1.58 $\pm$ 0.12
Progress Note GPT-4o 1.28 $\pm$ 0.03
o1-mini 1.46 $\pm$ 0.03
Gemini-1.5 1.02 $\pm$ 0.02
Llama3 8B 0.73 $\pm$ 0.04
Discharge Summary GPT-4o 1.30 $\pm$ 0.02
o1-mini 1.55 $\pm$ 0.03
Gemini-1.5 0.98 $\pm$ 0.02
Llama3 8B 0.60 $\pm$ 0.04
Table 1: Average facts per sentence by note type. Red indicates less facts than sentences meaning the LLM is likely failing to capture all facts.
3.1 RQ1 — Fact Decomposition
Fact Counts
For all note types, GPT-4o and o1-mini produce more facts per sentence than Gemini-1.5 and Llama3-8B (Table 1). For discharge summaries, o1-mini generates 1.55 facts per sentence compared to 0.98 for Gemini-1.5 and 0.60 for Llama3-8B.
Figure 4 illustrates the distribution of the coefficient of variation (CV) in generated facts across clinical note types, showing the variability in the number of facts generated per document by LLMs. Discharge summaries exhibited the highest median CV (0.45), followed by progress notes (0.37), procedure notes (0.27), and nursing notes (0.21). The lower variability observed in procedure and nursing notes is expected given shorter token lengths.
Figure 4: Distribution of coefficient of variation across documents. Higher values indicate higher variance in the number of facts per documents across LLMs.
Fact Similarities
Using Earth Mover’s Distance (EMD) with a cosine cost function, we found that GPT-4o and o1-mini produced the most similar fact decompositions, closely followed by Gemini-1.5, whose pairwise EMD values against these models were consistently within 0.02 units of each other (Figure 3). In contrast, Llama3’s decompositions were notably more divergent, with pairwise EMD scores roughly 0.03–0.05 units higher. Compared to the original source sentences, GPT-4o decompositions had the lowest mean EMD, while Llama3’s were highest. Differences were also influenced by note type: procedure notes showed minimal variation due to their structured format, nursing notes had slightly higher EMDs reflecting more dynamic content, and discharge summaries—being lengthy and complex—exhibited the highest EMD values.
Figure 5: Sentence atomicity by note type and LLM, shown as the normalized distribution of the estimated number of facts-per-sentence.
Figure 6: Overview of in-depth qualitative review with GPT-4o on the fact decomposition on twenty randomly-selected examples from FactEHR. We report percentage of correct, independent and atomic facts as annotated by medical expert.
Sentence Atomicity
Overall, o1-mini produced the most facts per sentence, followed closely by GPT-4o, both maintaining similar rates of zero-fact sentences (Figure 5). Gemini-1.5 had more zero-fact sentences than GPT-4o and o1-mini, but fewer than Llama3. Llama3 produced facts for only about 20% of sentences in progress notes and discharge summaries, and, except for nursing notes, generated fewer facts than sentences (Table 1).
Nursing notes have the highest proportion of sentences with three or more facts, reflecting their high information density. Progress notes, which tend to be longer and more narrative, have the lowest proportion of such sentences.
Qualitative Expert Evaluation
Expert reviewers assessed overall completeness of fact decompositions and the correctness, independence, and atomicity of individual facts. Completeness, defined as capturing all information in the source, is low: only 5% (1/20) of decompositions meet this criterion. For individual facts, reviewers judge whether each fact accurately reflected the source (correctness), stood on its own (independence), and represented a minimal unit of information (atomicity).
Across models, Llama3 had the highest rate of incorrect facts. Discharge summaries yield the least number of atomic facts, reflecting their complexity, while progress and procedure notes contain more granular and independent facts. Detailed results for GPT-4o are shown in Figure 6; additional results for other models and a summary of key metrics are reported in Appendix H and Table 2, respectively.
Correct Independent Atomic
GPT-4o 100 88 88
Gemini-1.5 96 94 81
Llama3-8B 45 95 74
Table 2: Overview of qualitative (fact-level) review on the fact decomposition for 20 notes. We report percentage of correct, independent and atomic facts as annotated by medical expert.
Procedure Note Nursing Note
$P*$ $R*$ $R*_{w}$ $F1*$ $F1*_{w}$ $P*$ $R*$ $R*_{w}$ $F1*$ $F1*_{w}$
GPT-4o 98.5 78.7 92.4 86.5 95.0 97.5 88.9 92.2 92.3 94.2
o1-mini 97.8 78.4 93.0 86.2 95.4 96.6 88.5 93.3 91.6 94.4
Gemini-1.5 95.9 64.2 85.7 77.0 91.7 93.2 77.3 83.7 84.4 88.3
Llama3-8B 84.2 49.4 73.7 62.0 82.3 84.1 56.6 69.9 65.1 75.6
Progress Note Discharge Summary
$P*$ $R*$ $R*_{w}$ $F1*$ $F1*_{w}$ $P*$ $R*$ $R*_{w}$ $F1*$ $F1*_{w}$
GPT-4o 97.4 78.1 85.2 85.8 89.9 97.0 79.0 86.9 86.8 91.1
o1-mini 95.2 79.7 85.7 86.6 89.7 94.8 81.1 88.2 86.7 90.0
Gemini-1.5 96.0 66.2 75.2 77.9 83.9 96.5 65.4 76.3 76.9 83.8
Llama3-8B 83.4 41.9 57.4 53.8 67.0 84.6 38.9 57.3 50.4 67.2
[…]
3.2 RQ2 — Entailment
GPT-4o achieve the best F1 scores on four of the five benchmark datasets, including the FactEHR development set (full set of results in Appendix G, Table 6), and was therefore selected as the entailment judge for all 987,266 entailment pairs in the FactEHR dataset. From these, we randomly sampled 1,036 pairs for manual annotation by seven clinical experts. On doubly annotated entailment pairs, Fleiss’ Kappa was 0.67, indicating substantial inter-rater agreement. Under this evaluation, GPT-4o achieves a recall of 0.96 and precision of 0.86 (Appendix G, Table 7), closely mirroring its performance on the FactEHR development set. Overall, these findings suggest that GPT-4o serves as a reasonable proxy for human judgment in entailment assessments.
3.3 RQ3 — Fact Verification
Using GPT-4o as the LLM entailment judge, we compute fact precision, fact recall (weighted and unweighted), and fact F1 (weighted and unweighted) for each fact decomposition across all note types (Table 3). GPT-4o and o1-mini emerge as top performers across all note types. For example, GPT-4o achieves a slightly higher fact F1 than o1-mini on procedure notes (86.5 vs. 86.2) and nursing notes (92.3 vs. 91.6), while o1-mini outperforms GPT-4o on progress notes (86.6 vs. 85.8). On discharge summaries, the two models are nearly tied (86.8 for GPT-4o vs. 86.7 for o1-mini). When considering the weighted fact F1 metric, o1-mini surpasses GPT-4o on procedure notes and nearly closes the gap on nursing notes. This suggests that the weighting scheme, which accounts for differences in fact counts per note, can influence which model is deemed superior.
Across all note types, fact precision remained consistently higher and less variable than fact recall—indicating that while the generated facts were generally accurate, many relevant facts were omitted, resulting in greater variability in recall and aligning with our qualitative findings (Appendix G, Figure 7).
4 Related Work
Because LLMs encode extensive clinical and biomedical knowledge (Singhal et al., 2023a; Agrawal et al., 2022; Munnangi et al., 2024), their evaluation is challenging. Models may not follow desired formats or may produce semantically equivalent but lexically divergent responses (Wadhwa et al., 2023; Agrawal et al., 2022). Prior work (Min et al., 2023; Xie et al., 2024; Tian et al., 2023) attempts to automatically assess factuality and completeness by decomposing text into atomic facts; for example, (Xie et al., 2024) use atomic facts for claim precision and recall calculations via entailment models. However, it remains unclear how effectively current models generate accurate, comprehensive atomic facts for clinical text.
While considerable progress has been made on atomic fact generation in general domains (Min et al., 2023) and biomedical literature (Wadden et al., 2020; Wright et al., 2022), there is little to no research focusing on clinical notes. Existing frameworks such as MAIRA-2 (Bannur et al., 2024) address fact-checking in radiology reports but not the broader range of clinical documents. Another, recent work has evaluated the role of context in fact decomposition, arguing for “molecular facts” that balance granularity with sufficient context. (Gunjal and Durrett, 2024).
5 Conclusion
In this study, we evaluate four LLMs’ ability to decompose over 2,000 real-world, EHR-derived clinical notes of four types into independent facts. We examine the length and similarity of the decompositions and employed GPT-4o as an entailment judge to assess their accuracy and completeness. There was up to a 2.6-fold difference in the number of facts generated by different LLMs. Clinician review confirmed that while the generated facts were generally accurate, important details were often omitted. To enable reproducibility, we release FactEHR, a dataset encompassing 8,665 fact decompositions and 987,266 entailment pairs from four note types and three institutions, including 1,036 pairs annotated by seven clinical experts. The results highlight the need to improve LLM-based fact decomposition methods for clinical documents to support the use of LLMs for healthcare use.
6 Limitations
First, our study focuses on four common note types, but many other specialized note types (e.g., ophthalmology) may yield even greater variability in fact decomposition, suggesting our error estimates are conservative. Second, we rely on a combination of open-source and proprietary models within a HIPAA-compliant environment, ensuring privacy but limiting reproducibility; we mitigate this by releasing our dataset. Thirdly, we did not control for prompt templates across models. Another limitation of this study is the lack of access to ground truth fact decompositions for clinical notes. This constraint forces us to rely on estimates and proxy measures, such as fact recall, to assess the completeness of model-generated fact decompositions. These metrics are inherently limited in their ability to capture the true completeness and accuracy of the decompositions. Lastly, since we examine only English documents, the generalizability of these findings to other languages remains uncertain.
Acknowledgements
This work is generously supported by the Mark and Debra Leslie endowment for AI in Healthcare. We would also like to thank the Clinical Excellence Research Center (CERC) at Stanford for their support. AS thanks the Paul and Daisy Soros Foundation and Knight Hennessy Scholars for their support.
Appendix A Dataset Details
A.1 Data Sources for Clinical Notes
For FactEHR, we sampled 2,168 clinical notes from three de-identified research datasets. We describe each of these datasets in detail:
MIMIC
MIMIC-III (Johnson et al., 2016) is a database of de-identified EHR comprising over 40k patients admitted to the intensive care unit of the Beth Israel Deaconess Medical Center between 2001 and 2012. MIMIC-III includes several note types, we sample progress notes, discharge summaries and nursing notes.
MIMIC-CXR
The MIMIC Chest X-ray Database v2.0.0 (Johnson et al., 2019) is a large publicly available dataset of chest radiographs in DICOM format with free-text radiology reports. The dataset contains 377,110 images corresponding to 227,835 radiographic studies performed at the Beth Israel Deaconess Medical Center in Boston, MA. We only use the free-text radiology reports for fact generation. We mapped the test splits from MIMIC-CXR with MIMIC-III and used a sub-sample of the split for our experiments.
CORAL
is a collection of a diverse set of 20 breast cancer and 20 pancreatic cancer patients from the University of California, San Francisco (UCSF) Information Commons, which contained patient data between 2012–2022, de-identified with Philter (Radhakrishnan et al., 2023). We use all the 172 progress notes for our experiments.
MedAlign
is collection of de-identified EHR data from Stanford Hospital and Lucile Packard Children’s Hospital. We sub-sample progress notes, nursing notes and discharge summaries from MedAlign (Fleming et al., 2023) which consists of 276 longitudinal EHRs, out of which we use 750 notes.
Note Type # of Notes
MIMIC MedAlign CORAL
Progress Note 250 250 172
Nursing Note 250 129 -
Discharge Summary 250 117 -
Procedure Note - 250 -
Radiology Note 500 - -
Total 1250 750 172
Table 4: Counts of notes by type and data source. Radiology notes are a subtype of procedure notes, and were sampled from both MIMIC-III and MIMIC-CXR.
Note Type # Tokens
Mean (SD) Min/Max
Progress Note 1510 (1199) [ 65, 6758 ]
Nursing Note 247 (168) [ 60, 977 ]
Discharge Summary 2253 (1041) [ 155, 7264 ]
Procedure Notes 339 (314) [ 62, 3047 ]
Table 5: Note token length summary statistics.
A.2 Entailment Data Sources
For entailment evaluation, we use existing NLI datasets. We describe them in detail, in the following paragraphs:
MultiNLI
(Williams et al., 2018b) The Multi-Genre Natural Language Inference corpus is a crowd-sourced collection of 433k sentence pairs annotated with textual entailment information. The labels include entailment, neutral and contradiction.
MedNLI
(Romanov and Shivade, 2018b) is a dataset annotated by healthcare professionals for the task of natural language inference (NLI) based on patient medical histories. The premise sentences in this dataset are sourced from the MIMIC-III (Johnson et al., 2016) database. They use Past Medical History section where annotators write alternative sentence which could be one of these entailment, neutral or a contradiction. The test split consists of 1422 such pairs along with the label.
Scitail
(Khot et al., 2018) The SCITAIL dataset is created for answering school-level science questions by converting each question and its correct answer into an assertive hypothesis (H). Relevant sentences are extracted from a large text corpus to serve as premises (P). Each premise-hypothesis pair is then annotated (via crowdsourcing) as entails or neutral. There are 2,126 pairs in the test split.
SNLI
(Bowman et al., 2015) The Stanford Natural Language Inference corpusis a collection of 570k human-written English sentence pairs manually labeled for balanced classification with the labels entailment, contradiction, and neutral. Their test split is 10,000 sentence pairs.
Appendix B Note Types
FactEHR  includes samples from four clinical note types. We outline the types and their clinical purpose below.
Procedure Note
Document medical procedures, including diagnostic and therapeutic interventions. Radiology reports, as a key example, detail findings from imaging studies like X-rays, CT scans, and MRIs, often comparing them to prior exams. Other examples include endoscopies, biopsies, and surgical procedures. These notes typically include the procedure performed, clinical indications, findings, and follow-up recommendations.
Nursing Note
Written by nurses to provide a systematic assessment of a patient’s condition across individual body systems (e.g., cardiovascular, neurologic) during a specific time frame. The focus of the note is primarily on the patient’s current status, with less emphasis on detailed planning for future care.
Progress Note
Written by physicians to summarize a patient’s medical status over the preceding day and outline the care plan for the following day. They include a review of significant events and relevant diagnostic tests conducted during that period. They include key events, diagnostic tests, a current patient exam, active medical issues.
Discharge Summary
Written by physicians to synthesize key medical information from a patient’s hospitalization, including clinical notes, diagnostic reports, and treatment plans. They provide a concise overview of the patient’s presentation, past medical history, key findings, future medical plans, and discharge medications for subsequent healthcare providers.
Appendix C Hyperparameters
For the fact decomposition task, we evaluated four models. GPT-4o, Gemini Flash 1.5, and LLaMA 3 8B were configured with a maximum token generation length of 4000 tokens, temperature of 0.01 and a top-p value of 0.9. Additionally, we included the o1-mini model, which used a temperature of 1, a top-p value of 1, and a maximum of 16,000 new tokens to allow for extra tokens generated in its chain of thought reasoning. At the time of our experiments o1-mini did not support other temperature or top-p settings. For the entailment task, we evaluated five models: GPT-4o, Gemini Flash 1.5, LLaMA 3 8B, LLaMA 3 70B, and GPT-4o-mini. These models were configured with a maximum token generation length of 25 tokens, with the same generation parameters as the fact decomposition task: a temperature of 0.01 and a top-p value of 0.9.
Appendix D Annotation Guidelines
The following instructions were provided to clinical annotators for evaluating entailment of premise-hypothesis pairs.
Annotation Instructions
  1. 1.
    You have two sheets: one containing the reference texts (labeled as 1 and 2) and one for annotating claims.
  2. 2.
    Start in the annotation sheet.
  3. 3.
    For each row, review the text in the claim column and determine if all the information in it can be completely inferred from the reference text specified in the reference_ID column.
  4. 4.
    Annotate as follows:
    • 1: If all of the information in the claim column can be completely inferred from the reference text.
    • 0: If any part of the claim contains information that is not present in the reference text.
  5. 5.
    If necessary, open the reference text sheet and annotation sheet in two separate windows for ease of comparison.
Examples
Example of a claim that can be completely inferred (annotation = 1): Reference text: Left pleural effusion increased from prior scan. Claim: There is a pleural effusion on the left side. Explanation: The claim is completely inferable from the reference text.
Example of a claim that can be completely inferred (annotation = 1): Reference text: FINDINGS: Nearly complete opacification of the left hemithorax is of increasing density since the recent prior CT. Claim: The results showed nearly complete opacification of the left hemithorax, which has increased in density since the last CT scan. Explanation: The claim is completely inferable from the reference text.
Example of a claim that cannot be completely inferred (annotation = 0): Reference text: Left pleural effusion increased from prior scan. Claim: The CT scan shows that the pleural effusion is increased from the prior scan. Explanation: There is no mention of a CT scan in the reference text.
Example of a claim that cannot be completely inferred (annotation = 0): Reference text: The patient has been evaluated for potential consolidation or pneumothorax. Multiple prior chest radiographs and a recent Chest CT were referred for comparison. Claim: Evaluate for evidence of consolidation or pneumothorax. COMPARISON: Multiple prior chest radiographs most recent on 1/15/22. Explanation: The reference text does not mention that the most recent scan was on 1/15/22.
Example of a claim that cannot be completely inferred (annotation = 0): Reference text: 1. The patient is intubated and has hypoxic respiratory failure. 2. A comparison is made with a prior study. 3. The cardiac size is normal. 4. The lines and tubes are in the standard position. 5. There are large right and moderate left pleural effusions that are grossly unchanged. 6. The right upper lobe opacity has improved consistent with improving atelectasis. 7. The pleural effusions are associated with atelectasis and are larger on the right side. 8. There is mild vascular congestion. Multiple prior chest radiographs and a recent Chest CT were referred for comparison. Claim: FINAL REPORT SINGLE FRONTAL VIEW OF THE CHEST REASON FOR EXAM: Intubated patient, hypoxic respiratory failure. Explanation: The claim contains several pieces of information that are not inferable from the reference text: 1) The fact that it is a final report, 2) It is a single frontal view, 3) The reason for the exam is hypoxic respiratory failure.
Appendix E Compute Environment
Experiments are performed in a HIPAA compliant APIs and local on-prem university compute environment using x Nvidia V100 GPUs. All compute environments supported HIPAA-compliant data protocols.
Appendix F Sentence Tokenization
Sentence tokenization errors resulting from formatting characters and tabular data are show in Table 8.
Appendix G Supplementary Results
G.1 Entailment model benchmarking
We report the performance of five LLMs (GPT-4o, GPT-4o-mini, Gemini-1.5-flash, Llama3-8b-Instruct, Llama3-70b-instruct) across five NLI datasets (FactEHR development set, MedNLI, SciTail, MultiNLI, and SNLI) in Table 6.
Dataset Model N Unparsable P R F1 Acc.
FactEHR GPT-4o 2468 0 90.8 95.7 93.2 90.0
Gemini-1.5 2369 0 89.1 96.5 92.6 89.0
Llama3-8 2468 0.36 86.2 44.7 58.8 55.2
Llama3-70 2468 0.166 88.18 90.12 89.14 84.24
4o-mini 2468 0 92.16 90.97 91.56 87.97
MedNLI GPT-4o 1422 0 94.7 76.0 84.3 90.6
Gemini-1.5 1422 0 93.8 67.5 78.5 87.7
Llama3-8 1422 0.001 65.2 86.7 74.5 80.2
Llama3-70 1422 0.004 80.93 87.76 84.21 89.03
4o-mini 1422 0 91.71 74.68 82.33 89.31
SciTail GPT-4o 2126 0 89.7 83.6 86.5 89.7
Gemini-1.5 2120 0 79.5 88.2 83.6 86.3
Llama3-8 2126 0.002 47.9 98.8 64.6 57.0
Llama3-70 2126 0.0117 70.49 93.35 80.33 81.89
4o-mini 2126 0 80 90.74 85.03 87.35
MultiNLI GPT-4o 9832 0 94.1 82.0 87.7 91.9
Gemini-1.5 9832 0 91.4 86.6 88.9 92.4
Llama3-8 9832 0.008 55.5 97.3 70.7 71.6
Llama3-70 9832 0.007 79.93 92.23 85.64 89.11
4o-mini 9832 0 90.87 86.49 88.62 92.18
SNLI GPT-4o 10000 0 93.1 87.0 90.0 93.4
Gemini-1.5 10000 0 93.0 86.1 89.4 93.1
Llama3-8 10000 0.0036 56.5 97.6 71.6 73.9
Llama3-70 10000 0.006 74.08 94.27 82.96 86.96
4o-mini 10000 0 91.77 86.1 88.85 92.72
Table 6: Entailment evaluation benchmarking results. Note that Gemini-1.5 flash failed to process some examples due to content moderation ("N" column). Responses whose JSON output were not parsable were coerced to "0". Bolded are the highest numbers (F1 and accuracy) of a model for each dataset. Our results show that GPT-4o performs well on four out of five benchmarking datasets, including the FactEHR validation set. Therefore, we selected GPT-4o as the final LLM entailment judge.
G.2 Fact Atomicity
We plot the total number of generated facts in figure 8, it varies significantly across LLMs and note types. The red line indicates the total sentence count per note type.
Figure 7: Comparison of fact precision and unweighted fact recall distributions across documents for each note type and fact decomposition model.
G.3 Fact Precision and Fact Recall
In Figure 7, we compare the distributions of fact precision and unweighted fact recall across documents for each note type and fact decomposition model. Across all note types and models, fact recall exhibits a much wider distribution than fact precision, indicating a higher variability in the model’s ability to generate comprehensive fact decompositions. This variability in recall suggests that while some documents achieve relatively high recall, many documents leave significant portions of the original information from the source notes unaccounted for missing facts in the fact decompositions.
In contrast, fact precision is generally higher and shows tighter distributions across all models and note types. This finding suggests that, for the facts that are generated, they are largely accurate and correctly entailed from the source document.
G.4 Entailment Model Validation
In table 7, we present our results of GPT-4o (our evaluator model) on the human evaluation entailment dataset.
Figure 8: The total number of generated facts varies significantly across LLMs and note types. The red line indicates the total sentence count per note type.
Sensitivity Specificity PPV NPV F1 Accuracy N
96.1 64.7 86.3 87.8 90.9 86.7 1036
Table 7: Performance of GPT-4o on the 1,036 entailment pairs labeled by clinical experts.
Note Type Junk Partial Examples
Discharge Summary 1.2% 3.6% ., Rel., 05:55, Trimethoprim/Sulfamethoxazole., SIRS/Sepsis, 2+, 05/21/2018, 5
Progress Note 0.7% 10.4% In, Unknown;, jpg], 6., 101 mEq/L, fevers., 24 hours, 28.6, 11/23/2015
Nursing Note 0.2% 3.7% Learning Preference, MAE’s., Pink/ruddy., 01/09/2016, PT., NPN, AFSF., bs:ronchi., 50
Procedures 9.0% 6.0% ___, _________, I., *, (Over), $\mid$, _______
Table 8: Example sentence tokenization errors with separate junk and partial error rates. Junk sentences are formatting characters (e.g. lines or tokens denoting split sections). Partial indicates an incomplete sentence that was incorrectly split during sentence parsing.
Appendix H Qualitative Evaluation
Expert reviewers assessed overall completeness of fact decompositions and the correctness, independence, and atomicity of individual facts. Detailed results for Llama3-8B and Gemini are shown in Figure 9 and Figure 10 respectively.
Figure 9: Overview of in-depth qualitative review with Llama3-8B on the fact decomposition on twenty randomly-selected examples from FactEHR. We report percentage of correct, independent and atomic facts as annotated by medical expert.
Figure 10: Overview of in-depth qualitative review with Gemini on the fact decomposition on twenty randomly-selected examples from FactEHR. We report percentage of correct, independent and atomic facts as annotated by medical expert.
**footnotetext: We did not have a comprehensive qualitative evaluation done for o1-mini due to time constraints.
Appendix I Cost
We report API costs for generating fact decomposition of our 2,168 notes and generating entailment labels for 987,266 entailment pairs in Appendix Table 9.
Task Model API cost (USD)
Entailment GPT-4o 10,162.72
Fact decomposition GPT-4o 264.26
Fact decomposition O1-mini 317.11
Fact decomposition Gemini 1.5 flash 5.20
Table 9: API costs for generating fact decomposition of our 2,168 notes and generating entailment labels for 987,266 entailment pairs.
Appendix J Prompts
We use two different prompts decribed as follows:
Fact Decomposition:
We adopt the prompt from Min et al. (2023), and include two in-context examples of fact decompositions of clinical notes written by a medical expert, illustrated in figure 11 The LLM is instructed to output independent facts as a delimited string. This approach resulted in fewer parsing errors than other methods.
Entailment Evaluation
We adopt the prompt from Xie et al. (2024), instructs the LLM to output a binary 0/1 indicator of entailment in JSON format (figure 12). This prompt was preferred over those requesting rationales or chain-of-thought reasoning due to its efficiency in minimizing output tokens, reducing inference times, and lowering computational costs, especially given the large number of premise-hypothesis pairs in the FactEHR dataset.
Figure 11: Prompt for fact decomposition.
Figure 12: Prompt for entailment evaluation with GPT-4o.